6. Assume the risk-free rate is 6% and the market risk premium is 6%. The stock of Physicians Care Network (PCN) has a beta of 1.5. The last dividend paid by PCN (D0) was $2 per share. What would PCN’s stock value be if the dividend were expected to grow at a constant rate of negative 5%.

Choice: $6.00

Choice: $9.50

Choice: $13.45

Choice: $17.60
Answer #1

Value of stock = expected dividend / (required return - constant growth rate)

expected dividend = current dividend * (1 + growth rate) = $2 * (1 + (-5%) = $1.90

required return = risk free rate + (beta * market risk premium) = 6% + (1.5 * 6%) = 15%

Value of stock = $1.90 / (15% - (-5%)) = $9.50